OpenStudy (anonymous):

find f ' (a) f(x)=x^-2

OpenStudy (anonymous):

f(a+h) - f(a) / h = (a+h)^-2 - a^-2 / h = ( 1/(a+h)^2 - 1/a^2 ) / h = a^2 - (a^2 + 2ah + h^2) --------------------- h(a^2+2ah+h^2)(a^2)

OpenStudy (anonymous):

f(a+h) - f(a) / h = (a+h)^-2 - a^-2 / h = ( 1/(a+h)^2 - 1/a^2 ) / h = a^2 - (a^2 + 2ah + h^2) --------------------- h(a^2+2ah+h^2)(a^2)

OpenStudy (anonymous):

just not sure where to go from here

OpenStudy (asnaseer):

expand the bracket in the numerator

OpenStudy (asnaseer):

the a^2 terms will cancel out in the numerator

OpenStudy (asnaseer):

then factorise the numerator and you should find and 'h' can cancel between numerator and denominator

OpenStudy (asnaseer):

finally take the limit as h tends to zero - which in this case you can just set h=0

OpenStudy (anonymous):

i got -2a --- a^4

OpenStudy (asnaseer):

that is correct - now simplify this

OpenStudy (anonymous):

-2/a^3

OpenStudy (asnaseer):

perfect! :)

OpenStudy (anonymous):

thanks!!!!

OpenStudy (asnaseer):

yw :)